OpenStudy (anonymous):

Solve the system of equations: what is the x-value for the solution? 2x + 7y = 3 and x = 1 - 4y A. -1 B. 5 C. 4 D. 3

OpenStudy (anonymous):

Explain please

OpenStudy (anonymous):

x = 5

OpenStudy (amistre64):

since we are given an equation that determines the value of x if y is known; we can substitute this into the x portion of the other equation

OpenStudy (anonymous):

y= (3-2x)/7 x=1-4(3-2x)/7

OpenStudy (amistre64):

x = 1 - 4y 2x + 7y = 3 --> 2(1 - 4y) +7y = 3 we solve this for y to determine x

OpenStudy (amistre64):

is one possible method .. ;)

OpenStudy (amistre64):

2-8y+7y = 3 2-y = 3 y = -1 ...................... x = 1-4y x = 1-4(-1) x = 1+4 = 5

OpenStudy (anonymous):

thank you soo much